# Magneto-Thomson and transverse Thomson effects in an interacting hadron gas in the presence of an external magnetic field

### v1

@@ III Results and Discussions
In this section, we discuss the results for higher-order thermoelectric coefficients of a hot and dense hadron gas in the presence of an external magnetic field. We also present a short description of the Thomson coefficient to show a comparative analysis between the magneto-Thomson coefficient and the Thomson coefficient. While the latter was estimated from the analysis of the Seebeck coefficient in our previous study [13], the results shown here are obtained from more complete and rigorous calculations. This allows for a qualitative understanding of how the presence of a magnetic field modifies the higher-order transport properties, particularly in relation to the temperature dependence and baryon chemical potential. We have considered four different versions of hadron resonance gas models, including a discrete particle spectrum having all the hadrons and their resonances up to the mass cutoff $\Lambda$ = 2.6 GeV [51]. To estimate the relaxation time of the hadrons, we follow the hard-scattering approximation. A uniform radius, $r_{m}$ = 0.2 fm for mesons and $r_{b,\bar{b}}$ = 0.62 fm for baryons and antibaryons throughout the calculations  [52]. For the VDWHRG model, we take van der Waals parameters as $a$ = 0.926 GeV fm3 and $b$ = (16/3)$\pi r^{3}$ [52, 49].
@@
Figure 2: Magneto-Thomson coefficient ($Th_{B}$) as a function of temperature at magnetic field $eB$ = 0.1 $m_{\pi}^{2}$ (upper panel) and $eB$ = 1.0 $m_{\pi}^{2}$ (lower panel) for baryon chemical potential at $\mu_{B}$ = 0.10 GeV (left), 0.30 GeV (middle), and 0.40 GeV (right).
@@
Fig. 1 represents the Thomson coefficient $Th$ as a function of temperature $T$ for three different baryon chemical potentials, i.e., at $\mu_{B}$ = 0.10, 0.30, and 0.40 GeV. The black solid line represents the Thomson coefficient obtained in the IHRG model, whereas the red dotted line represents the EVHRG model. The green and cyan dashed lines represent the same for the RMFHRG and VDWHRG models, respectively. It is observed that for the chosen range of temperature (0.08 - 0.18 GeV), $Th$ is negative for a hot and dense hadron gas. The $Th$ is universally decreasing as the temperature increases. However, the degree of decrease is different for different models. To define the net heat flow for a relativistic system, we need a conserved charge. For the study of the HRG system, the heat current corresponds to the net baryon current. It is observed that the baryons have a dominant contribution over the mesons to the Thomson coefficient. The mesons contribute to thermoelectricity only through the enthalpy of the system. In hadron gas, the entropy production for lighter mesons like pions, kaons, etc., is higher. Therefore, the enthalpy per baryon ($h$) exceeds the single-particle energy ($\omega_{i}$), which results in the negative values of $Th$ [13]. The positive (negative) values of the Thomson coefficient lead to the net absorption (release) of heat in the system. Therefore, a non-zero Thomson coefficient can significantly affect the cooling dynamics of the medium. The Thomson coefficient can be calculated as the product of temperature and the rate of change of the Seebeck coefficient with respect to temperature [13]; therefore, the higher the rate of change of the Seebeck coefficient with temperature, the higher the value of $Th$ in magnitude as a function of $T$. The coefficient $Th$ shows overlap at lower values of temperature nearly up to $\approx 0.14$ GeV for the case of $\mu_{B}$ = 0.10 GeV. As the value of $\mu_{B}$ increases, the models start deviating from each other even at low temperatures. Additionally, a noticeable flatness appears in the VDWHRG model in the temperature range around $T\approx$ 0.13-0.16 GeV for $\mu_{B}$ = 0.40 GeV, indicating a plateau-like behavior of the Thomson coefficient in this region. On the other hand, the constant decreasing rate of the Seebeck coefficient around the above-mentioned temperature range [13] leads to the constant slope. Hence, the $Th$ shows a non-monotonic behavior at high values of $\mu_{B}$.
@@
Figure 3: Transverse Thomson coefficient ($Th_{N}$) as a function of temperature at magnetic field $eB$ = 0.1 $m_{\pi}^{2}$ (upper panel) and $eB$ = 1.0 $m_{\pi}^{2}$ (lower panel) for baryon chemical potential at $\mu_{B}$ = 0.10 GeV (left), 0.30 GeV (middle), and 0.40 GeV (right).

### v2

@@ II.3 Higher-order thermoelectric coefficients
II.4 Landau quantization of charged hadrons in a magnetic field
In the presence of an external magnetic field, say along the z-axis, the dynamics of charged hadrons undergo a fundamental change due to Landau quantization. The transverse motion is no longer continuous but restricted to discrete Landau levels, while the longitudinal motion along the field direction remains continuous. This leads to a modification of the phase–space structure, such that the standard three-momentum integration is replaced by a summation over Landau levels and spin projections combined with an integration over the longitudinal momentum,
$\int\frac{d^{3}k_{i}}{(2\pi)^{3}}\longrightarrow\frac{|q_{i}|B}{2\pi}\sum_{l=0}^{\infty}\sum_{s_{z}}\int_{-\infty}^{\infty}\frac{dk_{z}}{2\pi},$ (80)
where $q_{i}$ is the electric charge of the $i$th hadron, $B$ denotes the external magnetic field strength, $l$ labels the Landau levels, and $s_{z}$ represents the spin projection. Here, we consider the hadrons having the spins 0, 1/2, and 1 [67]. For the charged hadrons, we also consider in the transverse plane, $k_{x}^{2}\approx k_{y}^{2}\approx(\frac{k_{x}^{2}+k_{y}^{2}}{2})=\frac{2l|{q_{i}}|B}{2}$ [12].
The corresponding single–particle dispersion relation also acquires a discrete structure,
$\omega_{c,i}^{z}(k_{z_{i}},l,s_{zi})=\sqrt{k_{z_{i}}^{2}+m_{i}^{2}+2(l+1/2-s_{zi})|q_{i}|B},$ (81)
which encodes the discrete orbital quantization due to the magnetic field together with the spin–magnetic field coupling. In the strong–field limit, the system approaches the lowest Landau level (LLL) approximation, where only the $l=0$ mode survives, effectively reducing the dimensionality of the momentum space.
These modifications have several important consequences. First, the density of states of charged hadrons is reshaped, altering the particle number distribution and bulk thermodynamic quantities such as pressure and energy density. Second, since scattering probabilities depend on the available phase space, the relaxation times are significantly modified in a magnetic background. As a result, all transport coefficients acquire non-trivial corrections that encode the anisotropy introduced by the magnetic field.
Our analysis incorporates Landau quantization explicitly in the hadronic phase, thereby ensuring that both the thermodynamic inputs and the transport coefficients reflect the microscopic changes induced by magnetic fields. This provides a more realistic and consistent description of the magnetic response of hadronic matter. To estimate the higher-order thermoelectric coefficients in the presence of a magnetic field, the defined integrals in Eqs. (56)-(59), for the charged hadrons, get modified as
$H_{1_{c,i}}^{z}$ $=g_{i}\frac{|q_{i}|B}{2\pi}\sum_{l=0}^{\infty}\sum_{s_{z}}\int_{-\infty}^{\infty}\frac{dk_{z}}{2\pi}$
$~~~~~~~~~~~~~~~~~~~~~\frac{\tau^{i}_{R}}{1+(\Omega_{c_{i}}\tau^{i}_{R})^{2}}\left(\frac{{l|{q_{i}}|B}}{\omega_{i}^{2}}\right)(-\frac{\partial f_{i}^{0}}{\partial\omega_{i}}),$ (82)
$H_{2_{c,i}}^{z}$ $=g_{i}\frac{|q_{i}|B}{2\pi}\sum_{l=0}^{\infty}\sum_{s_{z}}\int_{-\infty}^{\infty}\frac{dk_{z}}{2\pi}$
$~~~~~~~~~~~~~~~~~~~~~~\frac{\tau^{i}_{R}(\Omega_{c_{i}}\tau^{i}_{R})}{1+(\Omega_{c_{i}}\tau^{i}_{R})^{2}}\left(\frac{{l|{q_{i}}|B}}{\omega_{i}^{2}}\right)(-\frac{\partial f_{i}^{0}}{\partial\omega_{i}}),$ (83)
$H_{3_{c,i}}^{z}$ $=g_{i}\frac{|q_{i}|B}{2\pi}\sum_{l=0}^{\infty}\sum_{s_{z}}\int_{-\infty}^{\infty}\frac{dk_{z}}{2\pi}$
$~~~~~~~~~~~~~~~~~~~~~~\frac{\tau^{i}_{R}\omega_{i}}{1+(\Omega_{c_{i}}\tau^{i}_{R})^{2}}\left(\frac{{l|{q_{i}}|B}}{\omega_{i}^{2}}\right)(-\frac{\partial f_{i}^{0}}{\partial\omega_{i}}),$ (84)
$H_{4_{c,i}}^{z}$ $=g_{i}\frac{|q_{i}|B}{2\pi}\sum_{l=0}^{\infty}\sum_{s_{z}}\int_{-\infty}^{\infty}\frac{dk_{z}}{2\pi}$
$~~~~~~~~~~~~~~~~~~~~~~\frac{\tau^{i}_{R}\omega_{i}(\Omega_{c_{i}}\tau^{i}_{R})}{1+(\Omega_{c_{i}}\tau^{i}_{R})^{2}}\left(\frac{{l|{q_{i}}|B}}{\omega_{i}^{2}}\right)(-\frac{\partial f_{i}^{0}}{\partial\omega_{i}}).$ (85)
Hence, Landau quantization induces modifications to both leading- and higher-order thermoelectric coefficients.
II.5 Higher-order thermoelectric coefficients in the presence of a time-varying magnetic field
The Boltzmann transport equation for a single particle species, within the RTA and under the influence of an external electromagnetic field, is mentioned in Eq. 38. Here, we consider a time-dependent electromagnetic field parameterized as [68, 69]
$B=B_{0}\exp\left(-\frac{\tau}{\tau_{B}}\right),\qquad E=E_{0}\exp\left(-\frac{\tau}{\tau_{E}}\right),$ (86)
where $B_{0}$ and $E_{0}$ denote the initial magnitudes of the magnetic and electric fields, respectively, while $\tau_{B}$ and $\tau_{E}$ are their decay parameters. The proper time is represented by $\tau$. The exponential decay form of these fields follows naturally from Maxwell’s equations in a conducting medium, where finite electrical conductivity leads to temporal damping of the fields. Such forms typically emerge when Ohm’s law is incorporated into Maxwell’s equations. For simplicity, we focus on the temporal evolution of the fields in proper time $\tau$, neglecting spatial variations that are assumed to be smooth.
To solve Eq. (38), we introduce an ansatz that characterizes the deviation from equilibrium as [64, 15]
$\delta f_{i}=(\vec{k}_{i}\cdot\vec{\Omega})\frac{\partial f^{0}_{i}}{\partial\omega_{i}},$ (87)
where the vector $\vec{\Omega}$ encodes the influence of the temperature gradient, electric field, and magnetic field. It can be expressed as
$\vec{\Omega}$ $=~\alpha_{1}\vec{E}+\alpha_{2}\dot{\vec{E}}+\alpha_{3}\vec{\nabla}T+\alpha_{4}(\vec{\nabla}T\times\vec{B})+\alpha_{5}(\vec{\nabla}T\times\dot{\vec{B}})$
$+\alpha_{6}(\vec{E}\times\vec{B})+\alpha_{7}(\vec{E}\times\dot{\vec{B}})+\alpha_{8}(\dot{\vec{E}}\times\vec{B}),$ (88)
where $\alpha_{j}$ ($j=1,2,\ldots,8$) are unknown coefficients that determine the relative strength of each term in driving the system out of equilibrium. All contributions consistent with the CP symmetries of the electric and heat currents are retained, including the time-varying electromagnetic field, temperature gradients, and their cross-terms up to the leading order. In the case of a static electromagnetic field, the coefficients $\alpha_{2}$, $\alpha_{5}$, $\alpha_{7}$, and $\alpha_{8}$ vanish. Furthermore, when the chiral chemical potential is zero, the $\vec{B}$ and $\dot{\vec{B}}$ terms do not contribute to the current [68].
Using the coefficients $\alpha_{j}$, the simplified form of $\delta f_{i}$ can be written as (see Appendix of Ref. [12])
$\delta f_{i}$ $=\frac{-q_{i}\tau_{R}^{i}}{(1+\chi_{i})(1+\chi_{i}+\gamma_{i}^{2})}\Bigg[\Bigg((1+\chi_{i})+\frac{\chi_{i}(1+\chi_{i}-\gamma_{i}^{2})}{(1+\gamma_{i}^{2})}\Bigg)$
$)]\displaystyle(\vec{v}_{i}\cdot\vec{E})+\Bigg(\gamma_{i}(1+\chi_{i})+\frac{\chi_{i}\gamma_{i}(2+\chi_{i})}{(1+\gamma_{i}^{2})}\Bigg)\Big(\vec{v}_{i}\cdot(\vec{E}\times\hat{b}$$\Big)\Bigg$]
$+\frac{(\omega_{i}-\beta_{i}h)\tau_{R}^{i}}{T(1+\chi_{i})(1+\chi_{i}+\gamma_{i}^{2})}\Bigg[(1+\chi_{i})\Big(\vec{v}_{i}\cdot\vec{\nabla}T\Big)$
$+\gamma_{i}(1+\chi_{i})\Big(\vec{v}_{i}\cdot(\vec{\nabla}T\times\hat{b})\Big)\Bigg]\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (89)
For simplicity, we take $\tau_{E}=\tau_{B}$ and $\chi_{i}=\frac{\tau_{R}^{i}}{\tau_{B}}=\frac{\tau_{R}^{i}}{\tau_{E}}$. Here, $\gamma_{i}=\tau_{R}^{i}\Omega_{c_{i}}$, where $\Omega_{c_{i}}=\frac{q_{i}B}{\omega_{i}}$ is the cyclotron frequency of each hadron.
The induced electric current, obtained using $\delta f_{i}$, reads
$j^{l}$ $=\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}\vec{k}_{i}}{(2\pi)^{3}}v_{i}^{2}\frac{\tau_{R}^{i}}{(1+\chi_{i}+\gamma_{i}^{2})(1+\chi_{i})}$
$\times\Bigg[-q_{i}\Bigg((1+\chi_{i})+\frac{\chi_{i}(1+\chi_{i}-\gamma_{i}^{2})}{(1+\gamma_{i}^{2})}\Bigg)\delta^{lk}E^{k}$
$+\Bigg(\gamma_{i}(1+\chi_{i})+\frac{\chi_{i}\gamma_{i}(2+\chi_{i})}{(1+\gamma_{i}^{2})}\Bigg)\epsilon^{ljk}h^{j}E^{k}+\frac{(\omega_{i}-\beta_{i}h)}{T}$
$\Bigg((1+\chi_{i})\delta^{lk}\frac{\partial T}{\partial x^{k}}+\gamma_{i}(1+\chi_{i})\epsilon^{ljk}h^{j}\frac{\partial T}{\partial x^{k}}\Bigg)\Bigg]\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (90)
To simplify further, we choose the magnetic field along the $z$-axis, while the electric field and temperature gradient lie in the transverse (xxyy) plane. Under these conditions, the components of the electric current in the plane are expressed as
$j_{x}=$ $\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}\vec{k}_{i}}{(2\pi)^{3}}v_{i}^{2}\frac{\tau_{R}^{i}}{(1+\chi_{i}+\gamma_{i})(1+\chi_{i})}\times\Bigg[-q_{i}\Bigg((1+\chi_{i})+\frac{\chi_{i}(1+\chi_{i}-\gamma_{i}^{2})}{(1+\gamma_{i}^{2})}\Bigg)E^{x}+\Bigg(\gamma_{i}(1+\chi_{i})+\frac{\chi_{i}\gamma_{i}(2+\chi_{i})}{(1+\gamma_{i}^{2})}\Bigg)E^{y}$
$+\frac{(\omega_{i}-\beta_{i}h)}{T}\Bigg((1+\chi_{i})\frac{\partial T}{\partial x}+\gamma_{i}(1+\chi_{i})\frac{\partial T}{\partial y}\Bigg)\Bigg]\frac{\partial f^{0}_{i}}{\partial\omega_{i}},$ (91)
and
$j_{y}=$ $\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}\vec{k}_{i}}{(2\pi)^{3}}v_{i}^{2}\frac{\tau_{R}^{i}}{(1+\chi_{i}+\gamma_{i}^{2})(1+\chi_{i})}\times\Bigg[-q_{i}\Bigg((1+\chi_{i})+\frac{\chi_{i}(1+\chi_{i}-\gamma_{i}^{2})}{(1+\gamma_{i}^{2})}\Bigg)E^{y}+\Bigg(\gamma_{i}(1+\chi_{i})-\frac{\chi_{i}\gamma_{i}(2+\chi_{i})}{(1+\gamma_{i}^{2})}\Bigg)E^{x}$
$+\frac{(\omega_{i}-\beta_{i}h)}{T}\Bigg((1+\chi_{i})\frac{\partial T}{\partial y}-\gamma_{i}(1+\chi_{i})\frac{\partial T}{\partial x}\Bigg)\Bigg]\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (92)
Equations (91) and (92) can be compactly rewritten by defining the following integrals:
$H_{1_{i}}$ $=\frac{g_{i}}{3}\int\frac{d^{3}\vec{k}_{i}}{(2\pi)^{3}}\frac{\vec{k}_{i}^{2}}{\omega_{i}^{2}}~(-\frac{\partial f_{i}^{0}}{\partial\omega_{i}})\tau_{R}^{i}\times\frac{(1+\gamma_{i}^{2})+\chi_{i}(2+\chi_{i})}{(1+\chi_{i})(1+\gamma_{i}^{2})(1+\chi_{i}+\gamma_{i}^{2})},$ (93)
$H_{2_{i}}$ $=\frac{g_{i}}{3}\int\frac{d^{3}\vec{k}_{i}}{(2\pi)^{3}}\frac{\vec{k}_{i}^{2}}{\omega_{i}^{2}}~(-\frac{\partial f_{i}^{0}}{\partial\omega_{i}})\tau_{R}^{i}\times\gamma_{i}\frac{(1+\chi_{i})(1+\gamma_{i}^{2})+\chi_{i}(2+\chi_{i})}{(1+\chi_{i})(1+\gamma_{i}^{2})(1+\chi_{i}+\gamma_{i}^{2})},$ (94)
$H_{3_{i}}$ $=\frac{g_{i}}{3}\int\frac{d^{3}\vec{k}_{i}}{(2\pi)^{3}}\frac{\vec{k}_{i}^{2}}{\omega_{i}^{2}}(\omega_{i}-\beta_{i}h)~(-\frac{\partial f_{i}^{0}}{\partial\omega_{i}})\tau_{R}^{i}\times\frac{1}{(1+\chi_{i}+\gamma_{i}^{2})},$ (95)
$H_{4_{i}}$ $=\frac{g_{i}}{3}\int\frac{d^{3}\vec{k}_{i}}{(2\pi)^{3}}\frac{\vec{k}_{i}^{2}}{\omega_{i}^{2}}(\omega_{i}-\beta_{i}h)~(-\frac{\partial f_{i}^{0}}{\partial\omega_{i}})\tau_{R}^{i}\times\frac{\gamma_{i}}{(1+\chi_{i}+\gamma_{i}^{2})}.$ (96)
It is to be noted that these expressions are derived for the scenario of a time-varying magnetic field. For the case of a constant magnetic field ($\tau_{B}\rightarrow\infty$ or $\chi_{i}\rightarrow 0$ ), the expressions of components of electric currents in Eqs. (91)- (92) reduces to the Eqs. (54)- (55).
@@ III Results and Discussions
[…]
@@
Figure 2: Magneto-Thomson coefficient ($Th_{B}$) as a function of temperature at magnetic field eBeB = 0.1 $m_{\pi}^{2}$ (upper panel) and $eB$ = 1.0 $m_{\pi}^{2}$ (lower panel) for baryon chemical potential at $\mu_{B}$ = 0.10 GeV (left), 0.30 GeV (middle), and 0.40 GeV (right).
@@
Fig. 1 represents the Thomson coefficient $Th$ as a function of temperature $T$ for three different baryon chemical potentials, i.e., at $\mu_{B}$ = 0.10, 0.30, and 0.40 GeV. The black solid line represents the Thomson coefficient obtained in the IHRG model, whereas the red dotted line represents the EVHRG model. The green and cyan dashed lines represent the same for the RMFHRG and VDWHRG models, respectively. It is observed that for the chosen range of temperature (0.08 - 0.18 GeV), $Th$ is negative for a hot and dense hadron gas. The $Th$ is universally decreasing as the temperature increases. However, the degree of decrease is different for different models. To define the net heat flow for a relativistic system, we need a conserved charge. For the study of the HRG system, the heat current corresponds to the net baryon current. It is observed that the baryons have a dominant contribution over the mesons to the Thomson coefficient. The mesons contribute to thermoelectricity only through the enthalpy of the system. In hadron gas, the entropy production for lighter mesons like pions, kaons, etc., is higher. Therefore, the enthalpy per baryon ($h$) exceeds the single-particle energy ($\omega_{i}$), which results in the negative values of $Th$ [13]. The positive (negative) values of the Thomson coefficient lead to the net absorption (release) of heat in the system. Therefore, a non-zero Thomson coefficient can significantly affect the cooling dynamics of the medium. The Thomson coefficient can be calculated as the product of temperature and the rate of change of the Seebeck coefficient with respect to temperature [13]; therefore, the higher the rate of change of the Seebeck coefficient with temperature, the higher the value of $Th$ in magnitude as a function of $T$. The coefficient $Th$ shows overlap at lower values of temperature, nearly up to $\approx 0.14$ GeV for the case of $\mu_{B}$ = 0.10 GeV. As the value of $\mu_{B}$ increases, the models start deviating from each other even at low temperatures. Additionally, a noticeable flatness appears in the VDWHRG model in the temperature range around $T\approx$ 0.13-0.16 GeV for $\mu_{B}$ = 0.40 GeV, indicating a plateau-like behavior of the Thomson coefficient in this region. On the other hand, the constant decreasing rate of the Seebeck coefficient around the above-mentioned temperature range [13] leads to the constant slope. Hence, the $Th$ shows a non-monotonic behavior at high values of $\mu_{B}$. The qualitative differences observed among the various HRG model results can be attributed to the manner in which different interaction mechanisms modify the fundamental thermodynamic inputs that enter transport phenomena. In this context, it is instructive to consider the behavior of the electrical conductivity scaled by temperature, $\sigma_{\mathrm{el}}/T$, since electrical conductivity provides the leading contribution to thermoelectric transport coefficients and therefore serves as a useful baseline for understanding model-dependent trends. At higher temperatures, the number density of the medium increases rapidly, and the repulsive component present in interacting HRG models becomes increasingly significant, as discussed in the literature [57]. This repulsive interaction effectively reduces the number of particles per unit volume, which in turn enhances the relaxation time (79) relative to the IHRG scenario. As a consequence, a finite enhancement in the magnitude of $\sigma_{\mathrm{el}}/T$ and similarly in other transport coefficients is observed once hadronic interactions are incorporated. With increasing baryon chemical potential $\mu_{B}$, the baryon number density further increases, causing the deviations from the IHRG behavior to shift toward lower temperatures. Since the Thomson coefficients are constructed from the same underlying thermodynamic quantities, such as number density and enthalpy, as well as the same relaxation times, their model-dependent behavior naturally follows the same physical mechanism. This explains why differences among HRG models become more pronounced in regimes of higher temperature and baryon density.
@@
Figure 3: Transverse Thomson coefficient ($Th_{N}$) as a function of temperature at magnetic field eBeB = 0.1 $m_{\pi}^{2}$ (upper panel) and $eB$ = 1.0 $m_{\pi}^{2}$ (lower panel) for baryon chemical potential at $\mu_{B}$ = 0.10 GeV (left), 0.30 GeV (middle), and 0.40 GeV (right).